find the distance between S(5,1) and T(-3,3) using pythagorean theorem
without graphing the given points how can sides "a" and "b" be calculated???

Answer by checkley77(12844)

d^2=(3-1)^2+(5+3)^2
d^2=2^2+8^2
d^2=4+64
d^2=68
d=sqrt68
d=8.2462 is the distance between these two points.